# Compact stellar systems hosting an intermediate mass black hole: magnetohydrodynamic study of inflow-outflow dynamics

### v1

@@ 5.1 IMBH accretion fueled by stellar winds
Understanding accretion onto IMBHs in wind-fed systems is crucial for understanding their growth and observational signatures. In the compact WR stellar cluster scenario studied here, the winds create an unstructured inflow that affects accretion efficiency and detectability in X-rays and other wavelengths. As shown in Fig. 1, the interaction of these winds results in a turbulent, inhomogeneous flow with dense filaments and outflow channels. Furthermore, only about $10^{-5}$ of the injected mass reaches the black hole. This stark disparity highlights that only a tiny fraction of the available wind material is actually captured. Consequently, the black hole is unlikely to experience significant growth in mass from stellar winds alone. The accretion rates in our models are several orders of magnitude smaller than would be expected if the IMBH were fed by a dense gas disk or a continuous gas inflow (e.g., Poutanen et al., 2013). This implies that other processes—such as external gas infall from the larger environment, tidal disruptions of stars, or mergers with compact objects—would be needed for the IMBH to substantially increase its mass.
The accretion process is highly intermittent and variable, primarily driven by the orbital motions of the stars: periodic close WR-WR and WR-IMBH passages induce spikes in the inflow rate (see Fig. 10). These findings illustrate that wind-wind interactions dominate the gas dynamics, which is comprised almost entirely of unbound material. Instead, the resulting flow structure is largely outflow-dominated, which can further limit the amount of infalling gas. In summary, the wind-fed accretion in this environment is a fundamentally chaotic process, markedly different from the orderly disk accretion scenarios often considered for black hole growth.
@@
With such a meager feeding rate, the IMBH’s radiative output would be quite low. Assuming an advection-dominated (i.e. radiatively inefficient) accretion flow, we predict a quiescent X-ray luminosity of only $L_{X}\sim 10^{32}$ erg s-1, with short-lived bursts up to $\sim 10^{33}$ erg s-1 during quasi-periodic accretion spikes. These values are consistent with the observational upper limits for a putative IMBH in the IRS 13E cluster (Wang et al., 2020), and they are many orders of magnitude below the Eddington luminosity for a $3\times 10^{4}\,M_{\odot}$ black hole ($L_{\rm Edd}\sim 10^{42}\text{--}10^{43}$ erg s-1).
@@
The efficiency of radiative cooling—which depends on the wind’s metal content—strongly influences the morphology of the flow. In the high-metallicity case ($Z=0.4$, representative of WC8–9 type WR stars; Crowther 2007), cooling instabilities induce the formation of dense clumps via thin-shell shocks at wind–wind collision interfaces (e.g., Stevens et al., 1992; Calderón et al., 2025). But as noted above, these clumps do not substantially boost accretion. By contrast, in the lower-metallicity case the winds stay hotter and smoother, and the added thermal pressure further inhibits gas from falling inward.
@@
The specific angular momentum of the inflowing gas (Fig. 5) gives further insight into the structure and behavior of the accretion flow. Because the wind speeds are so high relative to the orbital speeds of the stars, most gas is injected into the domain with super-Keplerian angular momentum pointing in quasi-random directions. As a result, only the gas that is either directed very close to the black hole or gas that loses its angular momentum through wind-wind shocks can accrete. In the inner domain, this leads to a low angular momentum (both in terms of individual fluid elements and the net average) accretion flow without a coherent structure.
@@
Magnetic fields present in the stellar winds introduce additional physical effects. In the model with moderately magnetized winds ($\beta_{\rm w}=100$), we observe that magnetic pressure slows down the inflow and enhances turbulence. Magnetic dissipation converts some of the field energy into heat, raising the gas pressure and counteracting a portion of the radiative cooling. Although including magnetic fields does not dramatically change the global accretion rate in our simulations, it alters the geometry of outflows and suggests that even stronger wind magnetization could further inhibit accretion. $\beta$ profiles shown in Fig. 5 demonstrate that the combined thermal and ram pressure of the gas dominates the balance – evidenced by $\beta\gg 1$ at virtually all radii – meaning that the overall inflow–outflow structure is shaped primarily by hydrodynamic forces.
@@
Beyond accretion, the IMBH’s presence could influence the cluster’s overall dynamics. The kinetic energy imparted by the stellar winds and their interaction with the IMBH drive turbulence, potentially limiting the retention of cold gas in the cluster. Similar effects have been seen in AGN feedback studies involving observations, models or both (e.g., Zubovas & King 2012; King & Pounds 2015; Wylezalek & Zakamska 2016), where accretion flow-driven outflows can heat and expel surrounding material. This turbulence and gas expulsion may impact the long-term evolution of the cluster, affecting star formation and modifying the density profile, though these effects are not solely dependent on the IMBH’s presence. Future studies incorporating stellar evolution and external gas supply mechanisms will be crucial for understanding the full impact of IMBHs in such environments.
@@
In summary, wind-fed accretion onto an IMBH in a dense stellar cluster is a highly inefficient process. The high-velocity winds effectively starve the black hole, keeping its accretion luminosity extremely low. Consequently, the IMBH would remain practically invisible against the bright background of colliding-wind X-ray emission from the cluster (as we discuss later), even though its presence can still perturb the surrounding medium through gravity and outflows.
@@ 5.2 Comparison with other wind-fed accretion regimes
Wind-fed accretion has been most extensively studied in the context of Sgr A* – the supermassive black hole at the Galactic center – which is fed by the winds of nearby massive stars (e.g., Cuadra et al. 2008; Ressler et al. 2018, 2019a; Calderón et al. 2020b, 2025). Our work is the first to explore a similar wind-fed accretion scenario for an IMBH embedded in a compact star cluster. Comparing the two cases provides insight into how differences in black hole mass, stellar cluster size, wind properties, and cooling processes affect accretion dynamics and efficiency.
A key distinction between Sgr A* and our IMBH scenario is the physical scale and environment. Sgr A* ($M_{\bullet}\approx 4\times 10^{6}\,M_{\odot}$) lies in the Galactic nucleus, surrounded by a relatively extended cluster of OB-spectral type and WR stars spread over a region of a few tenths of a parsec. In contrast, our simulations consider a much smaller black hole ($M_{\bullet}=3\times 10^{4}\,M_{\odot}$) embedded in a very compact cluster (with radial extent on the order of 0.1 pc) of only a handful of WR stars. The stellar density in the IMBH’s immediate vicinity is, therefore, much higher in our scenario. On the other hand, the ratio between the gravitational capture radius of the IMBH and the scale of the cluster is comparable with the SMBH in the Galactic center.
Both Sgr A* and the simulated IMBH exhibit extremely low accretion efficiencies in their wind-fed modes. HD simulations of Sgr A*’s environment have shown that only a small fraction of the stellar wind material (of order $10^{-3}$) is captured by the SMBH, with the rest being blown away by outflows (Cuadra et al., 2008; Ressler et al., 2018). In our IMBH case, the captured fraction is even smaller, on the order of $10^{-5}$. In absolute terms, the time-averaged accretion rate for Sgr A* is estimated to be $\sim 10^{-9}\text{--}10^{-7}\,M_{\odot}\,\mathrm{yr}^{-1}$, whereas in our simulations it is $\sim 10^{-10}\,M_{\odot}$ yr-1. Despite the two-order-of-magnitude difference in black hole mass, the outcome is qualitatively similar: in both systems the vast majority of available wind mass is expelled and only a small fraction feeds the black hole.
The primary factor driving accretion inefficiency in both cases is the high wind velocity relative to the gravitational capture threshold. For an IMBH of mass $M_{\bullet}=3\times 10^{4}M_{\odot}$, the Bondi radius, which represents the scale at which gravitational influence becomes significant, is
@@
where $c_{\rm s}=\sqrt{\gamma{\rm k_{B}}T/\mu m_{\rm p}}$ is the sound speed of the gas. Assuming an adabiatic index of $\gamma=5/3$ and a resulting $\mathrm{k_{B}}T=3/16\,m_{\rm p}v_{\rm w}^{2}$ for strongly shocked gas, we can rewrite the Bondi radius in terms of wind speed $v_{\rm w}$ as
@@
In both systems, radiative cooling can lead to the formation of clumpy, filamentary structures in the accretion flow depending on the composition of the gas. In the Galactic center, these clumps can be bound and can therefore collect into an accretion disk and substantially raise the long-term accretion rate (e.g., Calderón et al., 2020b, 2025). In contrast, in our IMBH simulations, the dense clumps only momentarily increase local densities but are quickly disrupted or ejected, yielding no significant enhancement of accretion or disk formation.
Finally, while both observations of Sgr A* and our IMBH model display relatively low quiescent X-ray luminosities ($L_{X}\sim 10^{33}$ erg s-1 for Sgr A* and $\sim 10^{32}$ erg s-1 for the IMBH), they also show evidence of intermittent, episodic accretion events. Sgr A* exhibits frequent, daily flares in X-rays and near-infrared domains as well as at longer wavelengths (Baganoff et al., 2001; Porquet et al., 2003; Yusef-Zadeh et al., 2006; Genzel et al., 2010; Eckart et al., 2012, 2017; Witzel et al., 2021; von Fellenberg et al., 2025), thought to be triggered by sudden increases in accretion or magnetic reconnection events in the near-horizon flow that accelerate electrons to nonthermal energies (e.g., Dexter et al., 2020; Porth et al., 2021; Ripperda et al., 2022). In our simulations, we also observe short-lived accretion spikes when individual stars pass particularly close to each other and their wind collisions funnel gas inward. Although these bursts are much weaker relative to the “quiescent” X-ray emisison ($L_{\rm X}\lesssim 10^{32}\text{--}10^{33}$ erg s-1) than the observed Sgr A* flares (largest being $L_{\rm X}\lesssim 3-4\times 10^{35}$ erg s-1, 160 times the quiescent level Porquet et al. 2003), since our simulations do not capture the relevant scales or non-ideal physics important for such larger flares, we cannot rule them out. In fact, since both systems are in the radiatively inefficient regime with many qualitatively similar features in their larger scale flow, we posit that high magnitude flares originating from the near-horizon flow around the IMBH are likely (though on comparatively shorter timescales).
@@ 5.3 Observability
Our findings suggest that detecting an embedded IMBH in a WR stellar cluster poses a major observational challenge because the radiation from the accreting IMBH is overwhelmed by the emission from the surrounding stellar winds (as shown by our synthetic X-ray images in Fig. 6). Throughout our simulations, the vast majority of high-energy photons originate from wind–wind shock heating. This is true even when we account for the analytically predicted ADAF spectrum produced by the near-horizon flow (Figures 7 and 8), which results in an X-ray luminosity that is a factor of about $10^{2}\text{--}10^{3}$ smaller than that resulting from the wind-wind collisions. In absolute terms, the IMBH’s X-ray luminosity in our models (of order $10^{32}$ erg s-1) is extremely low – consistent with the non-detections and upper limits for IMBH candidates in similar environments (e.g., Chilingarian et al., 2018; Wang et al., 2020) – whereas the total diffuse X-ray luminosity of the cluster (from the stellar winds) is about $10^{34}\text{--}10^{35}$ erg s-1.
@@
The primary difficulty for observers, therefore, is one of contrast: disentangling the IMBH’s fainter radiative signal from the luminous stellar wind environment. In our scenario, the black hole would be effectively “lost in the glare” of the cluster’s collective emission. This emphasizes the need for sensitive, high-resolution observations to find an IMBH’s steady glow amid the cluster’s background. On the other hand, our simulations do not capture the physics and scales required for Sgr A*-like X-ray and infrared flares, meaning that there could still be strong transient signals. Evaluating this possibility is an important task for future work that includes such physics.
@@
Future X-ray and radio missions will provide enhanced sensitivity for detecting these objects. The proposed Lynx X-ray Observatory (Gaskin et al., 2019) will feature an angular resolution of $\sim 0.5$ arcseconds, comparable to Chandra X-ray Observatory, but with significantly improved sensitivity. This could allow for the detection of faint, persistent X-ray sources in dense stellar environments where IMBHs are suspected to reside. Given that the primary observational challenge is distinguishing the IMBH’s weak, point-like accretion signature from the surrounding diffuse stellar-wind emission, Lynx’s superior imaging and spectroscopic capabilities could be critical for resolving such sources. Similarly, the Advanced X-ray Imaging Satellite (AXIS) (Reynolds et al., 2023) is designed to provide high angular resolution X-ray imaging with a resolution of about 1.5 arcseconds across a wide field of view. Scheduled for launch in 2032, AXIS aims to study high-energy phenomena in the universe, including the detection of faint X-ray sources in dense stellar environments. Its rapid response capabilities to transient events and improved sensitivity make it a promising tool for identifying IMBHs. Looking further ahead, X-ray interferometry holds exciting potential for the future of high-resolution X-ray astronomy. By overcoming the diffraction limits of traditional telescopes, space-based X-ray interferometers could achieve microarcsecond angular resolution, making it possible to directly image IMBH environments and distinguish them from surrounding stellar activity (Uttley et al., 2021). Although still in an early conceptual phase, continued advancements in this technology could one day provide an unprecedented tool for confirming and characterizing IMBHs in dense stellar clusters.
Beyond X-rays, radio observations could offer complementary constraints. IMBHs accreting in a radiatively inefficient mode may exhibit faint synchrotron radio emission, analogous to low-luminosity AGN (Mezcua et al., 2018). Upcoming facilities such as the Square Kilometer Array (SKA) (Dewdney et al., 2009) could help probe this parameter space, particularly in nearby dense stellar clusters where IMBH candidates are suspected to reside. Furthermore, from the SED shown in Fig. 7 and after applying the extinction curve from Fritz et al. (2011), IRS 13E’s putative IMBH should exhibit radiative flux of 0.5 – 1.2 mJy at 19 $\mu$m. Assuming that the 25-80% FIR variability of Sgr A*’s peak thermal emission (von Fellenberg et al., 2018; Wielgus et al., 2022) is typical for low-luminosity accretion flows, this would result in $\sim$ 0.1 – 1 mJy mid-IR variability for the IRS 13E IMBH’s peak thermal emission, which should be detectable with JWST/MIRI (Wright et al., 2023).
@@
    Deep radio observations to constrain possible synchrotron emission from IMBHs accreting in a radiatively ineffecient mode.

### v2

@@ 5.1 IMBH accretion fueled by stellar winds
Understanding accretion onto IMBHs in wind-fed systems is crucial for understanding their growth and observational signatures. In the compact WR stellar cluster scenario studied here, the winds create an unstructured inflow that affects accretion efficiency and detectability in X-rays and other wavelengths. As shown in Fig. 1, the interaction of these winds results in a turbulent, inhomogeneous flow with dense filaments and outflow channels. Furthermore, only about $10^{-5}$ of the injected mass reaches the black hole. This stark disparity highlights that only a tiny fraction of the available wind material is captured. Consequently, the black hole is unlikely to experience significant growth in mass from stellar winds alone. The accretion rates in our models are several orders of magnitude smaller than would be expected if the IMBH fed a dense gas disk or a continuous gas inflow (e.g., Poutanen et al., 2013). This implies that other processes—such as external gas infall from the larger environment, tidal disruptions of stars, or mergers with compact objects—would be needed for the IMBH to increase its mass substantially.
The accretion process is highly intermittent and variable, primarily driven by the orbital motions of the stars: periodic close WR-WR and WR-IMBH passages induce spikes in the inflow rate (see Fig. 10). These findings illustrate that wind-wind interactions dominate the gas dynamics, which is comprised almost entirely of unbound material. Instead, the resulting flow structure is largely outflow-dominated, which can further limit the amount of infalling gas. In summary, wind-fed accretion in this environment is a fundamentally chaotic process markedly different from the orderly disk accretion scenarios often considered for black hole growth.
@@
The IMBH’s radiative output would be pretty low with such a meager feeding rate. Assuming an advection-dominated (radiatively inefficient) accretion flow, we predict a quiescent X-ray luminosity of only $L_{X}\sim 10^{32}$ erg s-1, with short-lived bursts up to $\sim 10^{33}$ erg s-1 during quasi-periodic accretion spikes, many orders of magnitude below the Eddington luminosity for a $3\times 10^{4}\,M_{\odot}$ black hole ($L_{\rm Edd}\sim 10^{42}\text{--}10^{43}$ erg s-1).
@@
The efficiency of radiative cooling—which depends on the wind’s metal content—strongly influences the morphology of the flow. In the high-metallicity case ($Z=0.4$, representative of WC8–9 type WR stars; Crowther 2007), cooling instabilities induce the formation of dense clumps via thin-shell shocks at wind–wind collision interfaces (e.g., Stevens et al., 1992; Calderón et al., 2025). Nevertheless, as noted above, these clumps do not substantially boost accretion. By contrast, in the lower-metallicity case, the winds stay hotter and smoother, and the added thermal pressure further inhibits gas from falling inward.
@@
The specific angular momentum of the inflowing gas (Fig. 5) gives further insight into the structure and behavior of the accretion flow. Because the wind speeds are so high relative to the orbital speeds of the stars, most gas is injected into the domain with super-Keplerian angular momentum pointing in quasi-random directions. As a result, only the gas that is either directed very close to the black hole or gas that loses its angular momentum through wind-wind shocks can accrete. In the inner domain, this leads to a low angular momentum (in terms of individual fluid elements and the net average) accretion flow without a coherent structure.
@@
Magnetic fields present in the stellar winds introduce additional physical effects. In the model with moderately magnetized winds ($\beta_{\rm w}=100$), we observe that magnetic pressure slows the inflow and enhances turbulence. Magnetic dissipation converts some field energy into heat, raising the gas pressure and counteracting a portion of the radiative cooling. Although including magnetic fields does not dramatically change the global accretion rate in our simulations, it alters the geometry of outflows. It suggests that even stronger wind magnetization could further inhibit accretion. $\beta$ profiles shown in Fig. 5 demonstrate that the combined thermal and ram pressure of the gas dominates the balance – evidenced by $\beta\gg 1$ at virtually all radii – meaning that the overall inflow–outflow structure is shaped primarily by hydrodynamic forces.
@@
Beyond accretion, the IMBH’s presence could influence the cluster’s overall dynamics. The kinetic energy imparted by the stellar winds and their interaction with the IMBH drives turbulence, potentially limiting the retention of cold gas in the cluster. Similar effects have been seen in AGN feedback studies involving observations, models, or both (e.g., Zubovas & King 2012; King & Pounds 2015; Wylezalek & Zakamska 2016), where accretion flow-driven outflows can heat and expel surrounding material. This turbulence and gas expulsion may impact the long-term evolution of the cluster, affecting star formation and modifying the density profile. However, these effects are not solely dependent on the presence of IMBH. Future studies incorporating stellar evolution and external gas supply mechanisms will be crucial for understanding the full impact of IMBHs in such environments.
@@
In summary, wind-fed accretion onto an IMBH in a dense stellar cluster is inefficient. The high-velocity winds effectively starve the black hole, keeping its accretion luminosity low. Consequently, the IMBH would remain practically invisible against the bright background of colliding-wind X-ray emission from the cluster (as discussed later), even though its presence can still perturb the surrounding medium through gravity and outflows.
@@ 5.2 Comparison with other wind-fed accretion regimes
Wind-fed accretion has been most extensively studied in the context of Sgr A* – the supermassive black hole at the Galactic center – which feeds the winds of nearby massive stars (e.g., Cuadra et al. 2008; Ressler et al. 2018, 2019a; Calderón et al. 2020b, 2025). Our work is the first to explore a similar wind-fed accretion scenario for an IMBH embedded in a compact star cluster. Comparing the two cases provides insight into how differences in black hole mass, stellar cluster size, wind properties, and cooling processes affect accretion dynamics and efficiency.
The physical scale and environment are key distinctions between Sgr A* and our IMBH scenario. Sgr A* ($M_{\bullet}\approx 4\times 10^{6}\,M_{\odot}$) lies in the Galactic nucleus, surrounded by a relatively extended cluster of OB-spectral type and WR stars spread over a region of a few tenths of a parsec. In contrast, our simulations consider a much smaller black hole ($M_{\bullet}=3\times 10^{4}\,M_{\odot}$) embedded in a very compact cluster (with radial extent on the order of 0.1 pc) of only a handful of WR stars. Therefore, the stellar density in the IMBH’s immediate vicinity is much higher in our scenario. On the other hand, the ratio between the gravitational capture radius of the IMBH and the scale of the cluster is comparable to that of the SMBH in the Galactic center.
Sgr A* and the simulated IMBH exhibit extremely low accretion efficiencies in their wind-fed modes. HD simulations of Sgr A*’s environment have shown that only a small fraction of the stellar wind material (of order $10^{-3}$) captured the SMBH, while the rest is blown away by outflows (Cuadra et al., 2008; Ressler et al., 2018). In our IMBH case, the captured fraction is even smaller, on the order of $10^{-5}$. In absolute terms, the time-averaged accretion rate for Sgr A* is estimated to be $\sim 10^{-9}\text{--}10^{-7}\,M_{\odot}\,\mathrm{yr}^{-1}$, whereas in our simulations it is $\sim 10^{-10}\,M_{\odot}$ yr-1. The outcome is qualitatively similar despite the two-order-of-magnitude difference in black hole mass. In both systems, most available wind mass is expelled, and only a small fraction feeds the black hole.
In both cases, the primary factor driving accretion inefficiency is the high wind velocity relative to the gravitational capture threshold. For an IMBH of mass $M_{\bullet}=3\times 10^{4}M_{\odot}$, the Bondi radius, which represents the scale at which gravitational influence becomes significant, is
@@
where $c_{\rm s}=\sqrt{\gamma{\rm k_{B}}T/\mu m_{\rm p}}$ is the sound speed of the gas. Assuming an adiabatic index of $\gamma=5/3$ and a resulting $\mathrm{k_{B}}T=3/16\,m_{\rm p}v_{\rm w}^{2}$ for strongly shocked gas, we can rewrite the Bondi radius in terms of wind speed $v_{\rm w}$ as
@@
In both systems, radiative cooling can lead to the formation of clumpy, filamentary structures in the accretion flow, depending on the gas composition. In the Galactic center, these clumps can be bound and can therefore collect into an accretion disk and substantially raise the long-term accretion rate (e.g., Calderón et al., 2020b, 2025). In contrast, in our IMBH simulations, the dense clumps only momentarily increase local densities but are quickly disrupted or ejected, yielding no significant enhancement of accretion or disk formation.
Finally, while both observations of Sgr A* and our IMBH model display relatively low quiescent X-ray luminosities ($L_{X}\sim 10^{33}$ erg s-1 for Sgr A* and $\sim 10^{32}$ erg s-1 for the IMBH), they also show evidence of intermittent, episodic accretion events. Sgr A* exhibits frequent, daily flares in X-rays and near-infrared domains as well as at longer wavelengths (Baganoff et al., 2001; Porquet et al., 2003; Yusef-Zadeh et al., 2006; Genzel et al., 2010; Eckart et al., 2012, 2017; Witzel et al., 2021; von Fellenberg et al., 2025), thought to be triggered by sudden increases in accretion or magnetic reconnection events in the near-horizon flow that accelerate electrons to nonthermal energies (e.g., Dexter et al., 2020; Porth et al., 2021; Ripperda et al., 2022). Our simulations also observe short-lived accretion spikes when individual stars pass particularly close to each other and their wind collisions funnel gas inward. However, these bursts are much weaker relative to the ”quiescent” X-ray emisison ($L_{\rm X}\lesssim 10^{32}\text{--}10^{33}$ erg s-1) than the observed Sgr A* flares (largest being $L_{\rm X}\lesssim 3-4\times 10^{35}$ erg s-1, 160 times the quiescent level Porquet et al. 2003), since our simulations do not capture the relevant scales or non-ideal physics important for such larger flares, we cannot rule them out. In fact, since both systems are in the radiatively inefficient regime with many qualitatively similar features in their larger-scale flow, we posit that high-magnitude flares originating from the near-horizon flow around the IMBH are likely (though on comparatively shorter timescales).
@@ 5.3 Observability
Our findings suggest that detecting an embedded IMBH in a WR stellar cluster poses a major observational challenge because the radiation from the accreting IMBH overwhelmed the emission from the surrounding stellar winds (as shown by our synthetic X-ray images in Fig. 6). Throughout our simulations, most high-energy photons originate from wind–wind shock heating. This is true even when we account for the analytically predicted ADAF spectrum produced by the near-horizon flow (Figures 7 and 8), which results in an X-ray luminosity that is a factor of about $10^{2}\text{--}10^{3}$ smaller than that resulting from the wind-wind collisions. In absolute terms, the IMBH’s X-ray luminosity in our models (of order $10^{32}$ erg s-1) is extremely low – consistent with the non-detections and upper limits for IMBH candidates in similar environments (e.g., Chilingarian et al., 2018; Wang et al., 2020) – whereas the total diffuse X-ray luminosity of the cluster (from the stellar winds) is about $10^{34}\text{--}10^{35}$ erg s-1.
@@
Therefore, the primary difficulty for observers is one of contrast: disentangling the IMBH’s fainter radiative signal from the luminous stellar wind environment. In our scenario, the black hole would be effectively ”lost in the glare” of the cluster’s collective emission, emphasizing the need for sensitive, high-resolution observations to find an IMBH’s steady glow amid the cluster’s background. On the other hand, our simulations do not capture the physics and scales required for Sgr A*-like X-ray and infrared flares, meaning that there could still be strong transient signals. Evaluating this possibility is an important task for future work that includes such physics.
@@
Future X-ray and radio missions will provide enhanced sensitivity for detecting these objects. The proposed Lynx X-ray Observatory (Gaskin et al., 2019) will feature an angular resolution of $\sim 0.5$ arcseconds, comparable to Chandra X-ray Observatory, but with significantly improved sensitivity. This could allow for detecting faint, persistent X-ray sources in dense stellar environments where IMBHs are suspected to reside. Given that the primary observational challenge is distinguishing the IMBH’s weak, point-like accretion signature from the surrounding diffuse stellar-wind emission, Lynx’s superior imaging and spectroscopic capabilities could be critical for resolving such sources. Similarly, the Advanced X-ray Imaging Satellite (AXIS) (Reynolds et al., 2023) is designed to provide high angular resolution X-ray imaging with a resolution of about 1.5 arcseconds across a wide field of view. Scheduled for launch in 2032, AXIS aims to study high-energy phenomena in the universe, including detecting faint X-ray sources in dense stellar environments. Its rapid response capabilities to transient events and improved sensitivity make it a promising tool for identifying IMBHs. Looking further ahead, X-ray interferometry holds exciting potential for the future of high-resolution X-ray astronomy. By overcoming the diffraction limits of traditional telescopes, space-based X-ray interferometers could achieve microarcsecond angular resolution, making it possible to directly image IMBH environments and distinguish them from surrounding stellar activity (Uttley et al., 2021). Although still in an early conceptual phase, continued advancements in this technology could one day provide an unprecedented tool for confirming and characterizing IMBHs in dense stellar clusters.
[…]
@@
    Deep radio observations to constrain possible synchrotron emission from IMBHs accreting in a radiatively inefficient mode.